OpenStudy (ray10):

Last one for tonight, Find a particular solution for \[\large \frac{d^{2}y}{dx}-4\frac{dy}{dx}+9y=4\cos(x)\]

OpenStudy (tkhunny):

Have you considered solving the Characteristic Equation, \(r^{2} - 4r + 9 = 0\)?

OpenStudy (ray10):

yes I did and I got \[\large m=2 \pm \sqrt5 i\] @tkhunny

OpenStudy (ray10):

well in this case \[\large r=2 \pm \sqrt5 i\]

OpenStudy (tkhunny):

Okay, since they are NOT Real, what's the plan?

OpenStudy (ray10):

as both r's are complex conjugates, I would have to find a corresponding solution, problem is, I don't know which solution :/

OpenStudy (tkhunny):

\(2 + \sqrt{5}i\implies e^{2t}\cos(\sqrt{5}t)\) \(2 - \sqrt{5}i\implies e^{2t}\sin(\sqrt{5}t)\)

OpenStudy (ray10):

okay, I see that would be taken from the table. so for a particular solution does that mean we can try, \[\large a*x*cos(kx)+b*x*sin(kx)\] ?

OpenStudy (tkhunny):

Well, you're not going to get \(cos(x)\) out of any of that. Are you SURE there IS a solution?

OpenStudy (ray10):

to be honest, I'm not 100% sure if there is one

OpenStudy (ray10):

the table always confuses me :S

OpenStudy (tkhunny):

If the only GENERAL solutions you are allowed are \(\cos(\sqrt{5}x)\;and\;\sin(\sqrt{5}x)\), I'm not seeing how we'll manage \(\cos(x)\).

OpenStudy (ray10):

I'm a little slow on that :( could you please clarify ?

OpenStudy (tkhunny):

\(\dfrac{d}{dx}\cos(\sqrt{5}x) = -\sqrt{5}\sin(\sqrt{5}x)\) \(\dfrac{d^{2}}{dx^{2}}\cos(\sqrt{5}x) = -5\cos(\sqrt{5}x)\) That pesky \(\sqrt{5}\) in the argument won't magically go away.

OpenStudy (ray10):

ohhh I see now, yes it will have to be incorporated through out the calculation :/

OpenStudy (ray10):

why did you second differentiate the \[\cos(\sqrt5 x)\] ?

OpenStudy (ray10):

oh never mind I see why

OpenStudy (ray10):

okay from having \[\large y'\] and \[\large y''\] where do we go from here?

OpenStudy (tkhunny):

Well, we know the general solution to the homogeneous equation. We have only to add sines and cosines to the general solution to find a solution to the nonhomogeneous equation. I get the homogeneous solution shown above, with some parameters to generalize and \(\dfrac{2}{5}\cos(x)-\dfrac{1}{5}\sin(x)\)

OpenStudy (ray10):

so that is the homogeneous solution to this question, from here what are the steps to work it out? I can try it

OpenStudy (tkhunny):

No, the solution to the homogeneous equation is the monstrosity with the \(\sqrt{5}\) wild. The sine and cosine just shown is one solution to the nonhomogeneous equation. However, this is not all. I used only sine and cosine. With two derivatives, we should also include x*sin(x) and x^2(sin(x)) along with x*cos(x) and x^2(cos(x)). After that, it is a very tedious simultaneous solution of 6 equations.

OpenStudy (ray10):

....6 equations :O :( I've been studying straight for seven hours, a question with 6 equations is going to hurt >.<

OpenStudy (ray10):

if I had the solution from this I could work it out and how I got to the solution from there, but there is no solutions for this >.<

OpenStudy (tkhunny):

Well, it may turn out that the last four I mentioned are all zero, so don't lose hope.

OpenStudy (ray10):

I shall try as much as I can, what are the equations for this large question?

OpenStudy (tkhunny):

I used \(f(x) = A\cos(x) + B\sin(x) + Cx\cos(x) + Dx\sin(x) + Ex^{2}\cos(x) + Fx^{2}\sin(x)\) Constructing \(\dfrac{d^{2}}{dx^{2}}f(x) - 4\dfrac{d}{dx}f(x) + 9f(x) = 4\cos(x)\), leads to the six equations. Like I said, it is tedious.

OpenStudy (ray10):

I see, hmm how do we get the six equations?

OpenStudy (tkhunny):

After expanding the massive expression on the left, group by the six function types: cos(x) sin(x) x*cos(x) x*sin(x) x^2 * cos(x) x^2 * sin(x) Each of these will have a coefficient made up of various combinations of A, B, C, D, E, and F. All but the coefficient on cos(x) are set to zero (0), since they do not appear int he particular solution. The one attached to cos(x) is 4.

OpenStudy (ray10):

so in use, the five equation below cox(x) are equal to zero

OpenStudy (ray10):

that means we are just dealing with cos(x) ?

OpenStudy (tkhunny):

No, the coefficients on the types are all zero. We are NOT just dealing with cos(x). We are dealing with anything that can result in a cos(x) up to the 2nd derivative. Briefly, if the particular solution is f(x) = Acos(x) + Bsin(x), we have f"(x) = -Acos(x) - Bsin(x) f'(x) = -Asin(x) + Bcos(x) f(x) = Acos(x) + Bsin(x) Giving: f"(x) - 4f'(x) + 9f(x) = (-Acos(x) - Bsin(x))-4(-Asin(x) + Bcos(x))+9(Acos(x) + Bsin(x)) = 4cos(x) Collecting: cos(x)[-A - 4B + 9A] + sin(x)[-B + 4A + 9B] = 4cos(x) For equality of the expressions, we must have: [-A - 4B + 9A] = 8A - 4B = 4 ==> 4A - 2B = 2 [-B + 4A + 9B] =4A + 8B = 0 ==> A + 2B = 0 This leads to 5A = 2 ==> A = 2/5 and finally B = -1/5, which is the solution I gave above.

OpenStudy (ray10):

oh my gosh, I totally understand that!!! :D

OpenStudy (tkhunny):

The trouble is wondering if that f(x) is sufficiently general. Do we need \(h(x) = f(x) + Cx\cos(x) + Dx\sin(x)\)? Do we need \(r(x) = h(x) + Ex^{2}\cos(x) + Fx^{2}\sin(x)\)? Note that the 1st and 2nd derivatives of h(x) and r(x) result in additional terms consisting of only cos(x). This may contribute to a more general solution.

OpenStudy (ray10):

ah and that's where I get down :/

OpenStudy (ray10):

and also, isn't \[\frac{2}{5}cos(x)-\frac{1}{5}sin(x)\] the final answer?? Because I just arrived at that step from my calculations

OpenStudy (tkhunny):

No worries. It's not massively difficult to demonstrate that you do not actually need these additional complications. IF we use just \(h(x) - f(x) = Cx\cos(x) + Dx\sin(x)\), because of symmetry we get the lovely result \(C = D = 0\). The same thing happens with \(r(x) - h(x) = Ex^{2}\cos(x) + Fx^{2}\sin(x)\). We get \(E = F = 0\) and I think we're done! We do NOT need to try anything with greater exponents, \(Gx^{3}\cos(x)\), for example. You might want to ponder why this is so. Now, go get some sleep!

OpenStudy (ray10):

oh my gosh, do you mean the final answer is \[\large \frac{2}{5}cos(x)-\frac{1}{5}sin(x)\] ???

OpenStudy (tkhunny):

I think you left of the solution to the homogeneous equation. Other than that, yes.

OpenStudy (ray10):

thank you soo very much @tkhunny what a great help you are :D yes I shall sleep now :D